OpenStudy (anonymous):

a 2 digit number is such that it tens digit is greater than it's units digit by four. the two-digit number is less than four times the product of its digits by 11..find the number

OpenStudy (anonymous):

Let \(n=10a+b\), where \(a,b\in\mathbb{Z}\), \(1\leq a\leq 9\), and \(0\leq b\leq 9\). We know that \(a=b+4\) and that \(n=10a+b=11b+40<44b^2+176b=4\cdot 11\cdot a\,b\).

OpenStudy (anonymous):

So \(0<44b^2+165b-40\), which is a simple quadratic...

OpenStudy (anonymous):

thanks